# On dense subalgebras of the singular ideal in groupoid C*-algebras

### v1

@@ 4. Ample Groupoids
Let $G$ be an ample groupoid. Then $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to the norm $\left\|\cdot\right\|_{I}$.
The norm $\left\|\cdot\right\|_{I}$ dominates the full norm $\left\|\cdot\right\|_{C^{*}(G)}$ and hence any $C^{*}$-norm. Therefore, $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
@@
Let $f\in J\cap\mathscr{C}_{c}(G)$. Take a compact set $K\subseteq G$ with ${\rm supp}^{\circ}(f)$ contained in its interior. Whenever $\dot{f}\in\mathscr{C}_{c}(G)$ satisfies ${\rm supp}^{\circ}(\dot{f})\subseteq K$ we have $\left\|\dot{f}\right\|_{I}\leq C_{K}\left\|\dot{f}\right\|_{\infty}$ where $\left\|\cdot\right\|_{\infty}$ denotes the supremum norm, and $C_{K}$ is a constant depending only on $K$. Therefore, it suffices to find, for each $\varepsilon>0$, a function $f^{\prime}\in J_{\mathbb{C}}$ satisfying ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$.
Fix $\varepsilon>0$. Take $x\in s(K)$, and let $g_{1},\dotsc,g_{m}\in G_{x}$ be all points in $G_{x}$ on which $f$ is non-zero. By Lemma 4.2, there exists a compact open neighbourhood $U_{x}\subseteq G^{(0)}$ of $x$, and compact open bisections $U_{1},\dotsc,U_{m}\subseteq K$ satisfying the following: $g_{i}\in U_{i}$ and $s(U_{i})=U_{x}$ for all $i=1,\dotsc,m$, and the function $f^{x}$ defined by
$f^{x}\coloneq\sum_{i=1}^{m}f(g_{i})\mathbbm{1}_{U_{i}}$
@@
lies in $J_{\mathbb{C}}$. Clearly ${\rm supp}^{\circ}(f^{x})\subseteq K$. We have $f|_{G_{x}}=f^{x}|_{G_{x}}$ and ${\rm supp}^{\circ}(f-f^{x})\subseteq K$, a compact set. Therefore, by Lemma 4.1, there exists a compact open neighbourhood $W_{x}\subseteq U_{x}$ of $x$ such that $\left|f(g)-f^{x}(g)\right|\leq\varepsilon$ whenever $s(g)\in W_{x}$. The sets $\{W_{x}\}_{x\in s(K)}$ form an open cover for the compact set $s(K)$, so select a finite subcover $W_{x_{1}},\dotsc,W_{x_{n}}$. Removing intersections if necessary, we may assume that the $W_{x_{j}}$ are pairwise disjoint (this is possible since the $W_{x_{i}}$ are contained in the Hausdorff space $G^{(0)}$, and are thus clopen in $G^{(0)}$). Define
$f^{\prime}\coloneq\sum_{j=1}^{n}(\mathbbm{1}_{W_{x_{j}}}\circ s)f^{x_{j}}.$
@@
Then $f^{\prime}\in J_{\mathbb{C}}$ satisfies ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f-f^{\prime}\right\|_{\infty}\leq\varepsilon$, as required. ∎
@@
Let $G$ be the groupoid arising from the action of a contracting self-similar group on a finite alphabet. Then $J_{\mathbb{C}}$ is dense in $J$.
@@
By Theorem 4.4 it suffices to check that $J\cap\mathscr{C}_{c}(G)$ is dense in $J$. By [23] and [15] the groupoid $G$ is second-countable and amenable. Moreover, [4, Corollary 7.13] implies that every element of $\tilde{G}^{(0)}$ is finite. The result follows by Corollary 3.12 (II). ∎

### v2

@@ 6.1. Building blocks for $J_{\mathbb{C}}$
Let $G$ be an ample groupoid that is covered by countably many open bisections. The algebraic singular ideal $J_{\mathbb{C}}$ is linearly spanned by elements of the form $f^{\Theta}$ from (6.1).
@@
Let $f\in J_{\mathbb{C}}$. Take a compact set $K\subseteq G$ with ${\rm supp}^{\circ}(f)\subseteq K$. Fix $x\in s(K)$, and let $g_{1},\dots,g_{n}$ be all elements of $G_{x}$ on which $f$ is non-zero. An argument identical to that in Lemma 5.4 yields compact open bisections $U_{1},\dots,U_{n}$ containing $g_{1},\dots,g_{n}$, with equal source, and such that
$f^{x}\coloneq\sum_{i=1}^{n}f(g_{i})\mathbbm{1}_{U_{i}}$
@@
is a finite sum of functions of the form $f^{\Theta}$ from (6.1). Let $U_{x}\coloneq s(U_{i})$ (note that this does not depend on $i$). In particular $x\in U_{x}$. We have $f|_{G_{x}}=f^{x}|_{G_{x}}$, so, by Lemma 5.1, $f-f^{x}$ is continuous at each $g\in G_{x}$. However, both $f$ and $f^{x}$ belong to $\mathbb{C}G$, and hence only take finitely many values. Therefore, by shrinking the $U_{i}$ if necessary, we can assume that $f|_{U_{x}}=f^{x}|_{U_{x}}$ (one can always shrink the compact open bisections in a tuple $\Theta$). The sets $\{U_{x}\}_{x\in s(K)}$ form an open cover for the compact set $s(K)$, so select a finite subcover $U_{x_{1}},\dotsc,U_{x_{m}}$. Removing intersections if necessary, we may assume that the $U_{x_{j}}$ are pairwise disjoint (this is possible since the $U_{x_{j}}$ are contained in the Hausdorff space $G^{(0)}$, and are thus clopen in $G^{(0)}$). Define
$f^{\prime}\coloneq\sum_{j=1}^{m}f^{x_{j}}.$
It is clear that $f^{\prime}$ is a finite sum of functions of the form $f^{\Theta}$ from (6.1). This completes the proof, since $f^{\prime}=f$ by construction. ∎
Remark 6.2.
The construction in this subsection generalises to ample groupoids that are not covered by countably many open bisections. After a modification analogous to that in Remark 5.6, one obtains, for any ample groupoid, and explicit spanning set for $J_{\mathbb{C}}$.
$\bullet$
6.2. Density of $J_{\mathbb{C}}$ in $J\cap\mathscr{C}_{c}(G)$
We prove that the algebraic singular ideal $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ for any ample groupoid – no countability assumptions are needed. In Corollary 6.4 this observation is applied to the class of groupoids arising from contracting self-similar groups.
Theorem 6.3.
Let $G$ be any ample groupoid. Then $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to the norm $\left\|\cdot\right\|_{I}$.
As stated earlier (see (5.4)), the norm $\left\|\cdot\right\|_{I}$ dominates the full norm $\left\|\cdot\right\|_{C^{*}(G)}$ and hence any $C^{*}$-norm. Therefore, $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
Proof.
Given $f\in J\cap\mathscr{C}_{c}(G)$, write $f=\sum_{j=1}^{m}f_{j}$ for some $f_{j}\in C_{c}(V_{j})$ and open bisections $V_{j}$, and let $H\subseteq G$ be the subgroupoid generated by the open bisections $V_{1},\dots,V_{m}$. Then $H$ is an open subgroupoid that is covered by countably many open bisections. By working in $H$ instead, we can assume that the groupoid $G$ is covered by countably many open bisections. Then, by Theorem 5.5, it suffices to assume that $f=f^{\Xi}$ for some tuple $\Xi=(x,h,b,U_{1},\dots,U_{n},\psi)$ as in Subsection 5.2, and where
$f^{\Xi}=\sum_{i=1}^{n}b(g_{i})(\psi\circ s|_{U_{i}}),$
@@ 6.2. Density of $J_{\mathbb{C}}$ in $J\cap\mathscr{C}_{c}(G)$
as in (5.3). Take a compact set $K\subseteq G$ containing each $U_{i}$. As in the proof of Theorem 5.5 it suffices to find, for each $\varepsilon>0$, a function $f^{\prime}\in J_{\mathbb{C}}$ satisfying ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f^{\Xi}-f^{\prime}\right\|_{\infty}\leq\varepsilon$.
Fix $\varepsilon>0$. Define $U\coloneq s(U_{i})$ (note that this does not depend on $i$). Since $U$ is totally disconnected, there exists a locally constant function $\varphi\in C_{c}(U)$ satisfying $\left\|\psi-\varphi\right\|_{\infty}\leq\frac{\varepsilon}{n\max_{i}(\left|b(g_{i})\right|)}$. Consider the function
$f^{\prime}\coloneq\sum_{i=1}^{n}b(g_{i})(\varphi\circ s|_{U_{i}}).$
It is clear that $f^{\prime}$ belongs to the Steinberg algebra $\mathbb{C}G$, and satisfies the conditions ${\rm supp}^{\circ}(f^{\prime})\subseteq K$ and $\left\|f^{\Xi}-f^{\prime}\right\|_{\infty}\leq\varepsilon$. Moreover, by Proposition 5.3, $f^{\prime}$ belongs to the singular ideal $J$, and hence also to $J_{\mathbb{C}}$. This completes the proof. ∎
@@
Let $G$ be the groupoid arising from a contracting self-similar group action. Then $J_{\mathbb{C}}$ is dense in $J$.
@@
By Theorem 6.3 it suffices to show that $J\cap\mathscr{C}_{c}(G)$ is dense in $J$. By [26, Corollary 2.19] (see also [29] and [15]) the groupoid $G$ is second-countable and amenable. Moreover, [5, Corollary 7.13] implies that every element of $\tilde{G}^{(0)}$ is finite. The result follows by Corollary 4.17 (II). ∎
Remark 6.5.
We would like to thank B. Kwaśniewski for pointing out to us that the previous result generalises to groupoids $G$ arising from contracting self-similar groupoid actions on finite graphs without sources. Such a groupoid is amenable by [26, Corollary 2.19], and it follows from either [23, Corollary 6.23] or [1, Proposition 9.1] that elements of $\tilde{G}^{(0)}$ are finite.